The scores of high school students on a national mathematics exam in SA were normally distributed with a mean of 86 and a standard deviation of 4. If there were 97680 students with scores higher than 91. how many students took the test? 394400 125000 247667 No correct answer provided 925000 105000


1
Expert's answer
2021-05-31T19:31:11-0400

Given,

"\mu=86\\\sigma=4\\"


So,

"z=\dfrac{91-\mu}{\sigma}=\dfrac{91-86}{4}=1.25"


So, "P(X>91)=P(z>1.25)"

From table:

"P(z>1.25)=0.1056"


and suppose 'n' students appear in the test.

So, according to question

"n\times 0.1056=97680\\\Rightarrow n=\dfrac{97680}{0.1056}=965000"


Hence, 965000 students appear in the test.